hey, how are you? I need help in this problem.
The length of a rectangle is 3ft more than twice its width, and the area of the rectangle is 35ft. Find the dimensions of the rectangle.
thanks for your time.
Juan

Answer by scott8148(6628)

let x=width, so 2x+3=length ... area=x(2x+3) ... 35=2x^2+3x

2x^2+3x-35=0 ... (2x-7)(x+5)=0 ... x=3.5 and x=-5 ... negative value not realistic